ISOMERIZATION BETWEEN CH_{2}ClI AND CH_{2}Cl-I IN CRYOGENIC MATRICES STUDIED ON ULTRA-FAST TIMESCALE

\underline{THOMAS J. PRESTON}, MAITREYA DUTTA, BRIAN J. ESSELMAN, MICHAEL A. SHALOSKI, ROBERT J. MCMAHON, and F. FLEMING CRIM, The University of Wisconsin-Madison Department of Chemistry, 1101 University Avenue, Madison, WI, 53705; AIMABLE KALUME, LISA GEORGE, and SCOTT A. REID, Department of Chemistry, Marquette University, Milwaukee, WI, 53233.

Photodissociation of species entrained in solid matrices potentially leads to reassociation of the newly formed fragments. After fixing CH_{2}ClI in various atomic and molecular matrices, we measure ultrafast transient absorptions to monitor the photolysis of the precursor and isomerization to form iso-CH_{2}ClI. We probe the two lowest energy electronic absorption features of CH_{2}Cl-I near 435 nm and 800 nm. Probing the low energy side of the 435-nm band interrogates the formation and subsequent cooling of the hot, newly formed products. We find that the recoiling fragments, CH_{2}Cl and I, lose large amounts of energy to the environment in the initial collision with the matrix cage, which leads to formation of the isomer.